# 2Huawei Noah’s Ark Lab, London, UK

### v1

@@ 4.2 Robot Experiments
Abbildung 9: Illustration of accuracy, robustness, and generalization evaluations. The accuracy evaluation requires the model to recognize the affordance of a single object and execute related task. The robustness evaluation involves accurately selecting a object in a cluttered scene to perform a specified affordance task. The generalization evaluation accesses if the model can reason about the graspable area of unseen objects.
4.2.1 Implementation Details
To evaluate the effectiveness of learned visual affordances, we deploy Aff-Grasp in a 7 DoF Kinova Gen3 robot arm. The arm is equipped with a Robotiq 2F-85 parallel jaw gripper, and a calibrated Azure Kinect RGB-D camera is mounted next to the robot to capture the scene of the workspace (see Figure LABEL:fig:setup). To enable open-vocabulary affordance recognition, we utilize CLIP text embeddings as the classifier, and discard the DFI to speed up inference time. Real-world experiments are conducted with 34 diverse objects (shown in Figure LABEL:fig:exp_objs) to evaluate three essential properties: accuracy, robustness, and generalization. We provide illustrations for these experiments in Figure 9, and introduce them in detail as follows:
  1. 1.
    Accuracy evaluation: Given a single seen object on the workspace, we evaluate whether the model can recognize correct affordances of the object and perform related affordance task. This evaluation is performed with 24 objects, each of which is repositioned 3 times during the experiment.
  2. 2.
    Robustness evaluation: Given multiple seen and unseen objects in a cluttered scene and an affordance task, we evaluate the model’s ability to identify which object should be selected to perform the specific task. This requires the model to make robust predictions in the presence of distractors. The evaluation is conducted across 7 affordance tasks. Each task is tested with 3 diverse objects, except for the handover task, which is tested with 6 objects, each possessing different functional affordances. Every object is repositioned 3 times during the experiments.
  3. 3.
    Generalization evaluation: Given novel objects not encountered during training, we evaluate if the model can still recognize the correct graspable areas. This evaluation assess if the model can generalize the graspable affordance prediction to novel objects, which is a crucial factor in robotic manipulation. It is conducted with 7 novel objects, each repositioned 5 times.
Success rate is adopted as metric and reported from three aspects: correct affordance prediction, successful grasp, and successful interaction. For experiments in cluttered scenes, we assume that only one object is available to complete the target task. We do not perform manipulation policy learning, as it is beyond the focus of this work. Instead, we design motion primitives for each affordance and assume that the operating direction of the tool is known.
Tabelle 3: Success rates for accuracy evaluation.
@@ 4.2.1 Implementation Details
LOCATE locate 42/72 (58.3%) 33/72 (45.8%) n/a
Robo-ABC robo-abc 62/72 (86.1%) 44/72 (61.1%) n/a
@@
Tabelle 4: Success rates for robustness evaluation.
Cut Stir Scoop Screw Pour Stick Handover Total
Correct Affordance 8/9 9/9 9/9 8/9 9/9 9/9 17/18 69/72 (95.8%)
Successful Grasp 7/9 6/9 7/9 6/9 7/9 6/9 13/18 53/72 (73.6%)
Successful Interaction 7/9 6/9 5/9 5/9 6/9 5/9 11/18 46/72 (63.9%)
Tabelle 5: Success rates for generalization evaluation and inference time for affordance prediction components.
@@
LOCATE locate 20/35 (57.1%) 15/35 (42.9%) 0.0047
Robo-ABC robo-abc 24/35 (68.6%) 21/35 (60.0%) 12.92
@@
4.2.2 Comparison Methods
We choose two relevant methods, LOCATE locate and Robo-ABC robo-abc , that learn affordances in a similar manner for comparison.
  • LOCATE: It is a state-of-the-art affordance grounding model that learns affordances from human-object interaction images using action labels as weak supervision. The method builds on DINO-ViT to identify object parts by clustering visual features from interaction regions of exocentric images, and then transfers the discovered parts to egocentric images for affordance grounding.
  • Robo-ABC: It extracts object images and contact points from egocentric videos and store these as an affordance memory. During inference, it first retrieves the most similar objects to the target and then utilizes semantic correspondence from the diffusion model to map the contact point to the current object.
Abbildung 10: Qualitative comparison of affordance prediction and final grasp pose for 3D point clouds in the cluttered scene. LOCATE fails to identify related objects for desired tasks, whereas Aff-Grasp can select the correct object with accurate affordance segmentation and is not affected by cluttered scenes.
Abbildung 11: Qualitative comparison of graspable affordance predictions on seen and unseen object categories.
4.2.3 Quantitative and Qualitative Comparisons
@@
The results for the accuracy evaluation are shown in Table 3. It is clear that success rates of Aff-Grasp significantly exceed those of its competitors, with 11.1% higher affordance prediction rate and 19.5% higher in successful grasping compared to Robo-ABC. Also, it is worth mentioning that our success rate for affordance prediction is measured on both graspable and functional affordances, whereas other two methods are measured solely on the graspable affordance. We observe that LOCATE struggles to make accurate predictions in real-world scenarios. Robo-ABC has relatively accurate affordance predictions, but it generates grasp proposals based on point correspondences, which do not represent the most confident grasp. Consequently, even though Robo-ABC frequently makes correct affordance predictions, 29% of its proposed grasping points do not lead to successful grasps.
In addition, we note that Aff-Grasp is capable of recognizing the correct affordance in cluttered scenes. As presented in Table 4 for the robustness evaluation, Aff-Grasp achieves a high success rate in affordance prediction, accurately predicting affordances 95% of the time, even in the presence of multiple seen and unseen objects acting as distractors. Table 5 reports results from the generalization evaluation and the inference time for the affordance prediction component of the models. It can be observed that our method is efficient and significantly more accurate in predicting the correct graspable areas for unseen objects, leading to a much higher success rate in grasping. While LOCATE also has a fast inference speed, it fails to accurately infer graspable affordances. In contrast, Robo-ABC’s performance on unseen objects is considerably reduced and suffers from a much longer inference time. Although it does not require additional training, its retrieval and correspondence mapping processes are quite time-consuming and computationally expensive, making it less suitable for real-world applications. In Figure 12, we show success rates of individual classes for accuracy and generalization evaluations. It is evident that our results are accurate and stable over all categories, while the results of LOCATE and Robo-ABC show frequent fluctuations.
[…]
Abbildung 12: Success rates of correct affordance predictions for each individual object from the accuracy and generalization evaluations. The total numbers of trials are 9 and 5, respectively.
Abbildung 13: The Aff-Grasp framework can perform the handover task by generating grasp poses within the functional parts of objects, and orienting the graspable parts towards the human hand. Green indicates all potential grasps, while red marks the final selected grasp.
@@
Qualitative comparisons are illustrated in Figure 10 and 11. In Figure 10, we display affordance segmentation maps and grasp poses for point clouds in a cluttered scene, where both seen and unseen objects serve as interferences (robustness evaluation). It is noted that LOCATE is unable to localize the correct object to execute the specified affordance task, while our model successfully identifies the matching object and predicts accurate segmentation maps. In Figure 11, we present raw predictions of graspable affordance from each model for seen and unseen objects. It can be observed that LOCATE often produces incomplete and wrong predictions. Robo-ABC occasionally makes prediction within the right object part area, but also produces high activation for the background or the entire object. In comparison, Aff-Grasp consistently makes precise segmentation predictions for both seen and unseen objects and is not affected by the background. Furthermore, we display affordance and grasp pose predictions for the handover task in Figure 13. When the robot is asked to pass something to the subject for a task, Aff-Grasp generates grasp proposals based on the functional affordance mask and directs graspable parts towards the subject’s hand.
@@ 4.3 Ablation Study
Tabelle 6: Ablation results of embedder, loss functions, classifiers, and proposed modules. The baseline model is a DeiT III model with a linear layer and binary cross entropy loss. “w/o bg” means that there is no background classifier. “DFI-training only” denotes that the DFI is only used during training, and discarded at inference.
@@
To explore the impact of each component in our model, we perform ablation experiments on the embedder, loss function, classifiers, designed modules, and hyper-parameters. The ablation results are summarized in Table 6. We first set up a baseline model, which employs a frozen DeiT III deit3 backbone that is fully supervised on ImageNet-1k. We then add a simple linear layer for patch-wise classification and utilize binary cross-entropy as the loss function. Based on this baseline, we first explore the impact of the embedder and loss functions. We find that a larger feature map followed by an embedder is beneficial, and the combination of focal loss and dice loss also brings improvements. Then, we analyze the results under different classification schemes, including the linear layer, cosine similarity, and whether to learn a background classifier. It is clear that implicit background prediction leads to better performance. Given the large gap between training and evaluation data, learning a background classifier can easily result in overfitting. Also, employing cosine similarity as the classifier can better utilize the inherent features of DINOv2, producing better results than a linear classifier. Lastly, we investigate the influence of DFI and LoRA. Notably, DFI improves performance significantly by a large margin, with 7.96% and 6.35% increases in mIoU and F1 score.
Abbildung 14: Visualization of qualitative improvements with DFI and LoRA.
In particular, DFI can also be used solely in training and discarded at inference, thereby improving results without extra computational cost. Additionally, integrating LoRA layers to fine-tune the foundation features is also helpful, leading to a 3.96% improvement in mIoU with marginal additional parameters. In Figure 14, we show the qualitative ablation results to visually examine the effects of DFI and LoRA. The segmentation results indicate that DFI is particularly effective at locating tiny and slender parts, while LoRA further enhances performance with refined boundaries and more complete segmentation maps.
Tabelle 7: Ablation study on different depth representations in DFI.
Depth map mIoU F1 Accuracy
Color depth (jet) 61.82 76.29 78.34
Color depth (inferno) 62.38 76.57 77.31
Color depth (viridis) 63.92 77.81 78.95
Grayscale depth 64.66 78.35 79.74
To further understand the effectiveness of DFI module, we perform experiments using different depth maps as input. We observe that DFI is more effective with depth representations that have low contrast. As listed in Table 7, the jet colormap, known for high-contrast visual effect, yields the worst results in DFI. In comparison, the less expressive grayscale depth achieves the best performance among other colored counterparts. We speculate that grayscale input focuses more on the geometric information, whereas color depth may introduce noise to some extent. In Table 8, we show the impact of different classification embeddings. The learnable embeddings yield the best results, but lose the ability to reason about unseen affordances. While performance degrades slightly when using CLIP text embeddings as classifiers, this approach retains the ability for open-vocabulary affordance segmentation. Therefore, we use learnable embeddings for vision evaluation and CLIP-L/14 text embeddings for robot experiments.
Tabelle 8: Ablation study on different classification embeddings: learnable or CLIP text embeddings.
Embeddings mIoU F1 Accuracy
CLIP-B/32 66.47 79.70 79.31
CLIP-B/16 66.04 79.37 79.15
CLIP-L/14 66.91 80.02 81.09
Learnable embeds 68.62 81.09 83.51
Finally, we conduct experiments with different hyer-parameter settings, focusing on the threshold $\tau$ for background classification and the weighting factor $\alpha$ in the loss function. As presented in Figure 15, the model obtains the highest performance in mIoU and F1-score with a weighting factor $\alpha$ of 1. For the background classification threshold $\tau$, a smaller value leads to higher accuracy, as only confident predictions are counted as foreground. In this case, only mIoU and F1 score can truly reflect the performance. We thus choose 0.8 as the default threshold.
@@
Abbildung 15: Ablation study on hyper-parameters.

### v2

@@ 4.2 Robot Experiments
Figure 6: Illustration of accuracy, robustness, and generalization evaluations. (a) The accuracy evaluation requires the model to recognize the affordance of a single object and execute related tasks. (b) The robustness evaluation involves accurately selecting an object in a cluttered scene to perform a specified affordance task. (c) The generalization evaluation succeeds if the model can reason about the graspable area of unseen objects.
Experiment Setup and Comparison Methods. Real-world experiments are conducted to evaluate three essential properties: accuracy, robustness, and generalization. Illustrations for these experiments are shown in Fig. 6. For the comparison methods, since this work focuses on visual affordance learning (i.e., identifying appropriate graspable and functional regions) rather than grasp success, manipulation models are not directly comparable. Instead, we select two relevant methods, LOCATE [23] and Robo-ABC [19], that learn affordances in a similar manner—from egocentric images and videos, respectively.
@@
LOCATE [23] 42/72 (58.3%) 33/72 (45.8%) n/a
Robo-ABC [19] 62/72 (86.1%) 44/72 (61.1%) n/a
@@
Table 2: Success rates for accuracy evaluation. “n/a” denotes inapplicability to tool-object interactions, as LOCATE’s predictions often overlap, and Robo-ABC cannot infer functional areas.
@@
LOCATE [23] 20/35 (57.1%) 15/35 (42.9%) 0.0047
Robo-ABC [19] 24/35 (68.6%) 21/35 (60.0%) 12.92
@@
Table 3: Success rates for generalization evaluation and inference time for affordance prediction components.
@@
Quantitative and Qualitative Comparisons. The results for the accuracy evaluation are shown in Tab. 2. It is clear that Aff-Grasp outperforms its competitors significantly, achieving an 11.1% higher affordance prediction rate and a 19.5% increase in successful grasping compared to Robo-ABC. Also, it is worth mentioning that our success rate for affordance prediction is measured on both graspable and functional affordances, whereas other two methods are measured solely on the graspable affordance. We observe that LOCATE struggles to make accurate predictions in real-world scenarios. Robo-ABC has relatively accurate affordance predictions, but it generates grasp proposals based on point correspondences, which do not represent the most confident grasp. Consequently, even though Robo-ABC frequently makes correct affordance predictions, 29% of its proposed grasping points do not lead to successful grasps. In addition, we note that Aff-Grasp is capable of recognizing the correct affordance in cluttered scenes (robustness evaluation). It achieves a high success rate in affordance prediction, accurately predicting affordances 95% of the time, even in the presence of multiple seen and unseen objects acting as distractors (results and visual examples are provided in the supplementary). Table 3 reports results from the generalization evaluation and the inference time for the affordance prediction component of the models. It can be observed that our method is efficient and significantly more accurate in predicting the correct graspable areas for unseen objects, leading to a much higher success rate in grasping.
@@
In Fig. 7, we present raw predictions of graspable affordance from each model for seen and unseen objects. It can be observed that LOCATE often produces incomplete and wrong predictions. Robo-ABC occasionally makes predictions within the right object part area, but also produces high activation for the background or the entire object. In comparison, Aff-Grasp consistently makes precise segmentation predictions for both seen and unseen objects and is not affected by the background.
Figure 7: Qualitative comparison of graspable affordance predictions on seen and unseen object categories.
@@ 4.3 Ablation Study
Table 4: Ablation results of embedder, loss functions, classifiers, and proposed modules. The baseline model is a DeiT III model with a linear layer and binary cross entropy loss. “w/o bg” means that there is no background classifier. “DFI-training only” denotes that the DFI is used during training, and discarded at inference.
Inference #Params (M) GFLOPs Time (ms)
@@
[…]
Table 5: Ablation study on inference efficiency of DFI.
To explore the impact of each component in our model, we perform ablation experiments on the embedder, loss function, classifiers, and designed modules. The ablation results are summarized in Tab. 4. We first set up a baseline model, which employs a frozen DeiT III [51] backbone that is fully supervised on ImageNet. We then add a simple linear layer for patch-wise classification and utilize binary cross-entropy as the loss function. Based on this baseline, we first explore the impact of the embedder and loss functions. We find that a larger feature map followed by an embedder is beneficial, and the combination of focal loss and dice loss also brings improvements. Then, we analyze the results under different classification schemes, including the linear layer, cosine similarity, and whether to learn a background classifier. It is clear that implicit background prediction leads to better performance. Given the large gap between training and evaluation data, learning a background classifier can easily result in overfitting. Also, employing cosine similarity as the classifier can better utilize the inherent features of DINOv2, producing better results than a linear classifier. Lastly, we investigate the influence of DFI and LoRA. Notably, DFI improves performance significantly by a large margin, with 7.96% and 6.35% increases in mIoU and F1 scores.
Figure 8: Qualitative improvements with DFI and LoRA.
In particular, DFI can also be used solely in training and discarded at inference, improving results without extra computational cost. The impact of this operation is analyzed in Tab. 5, showing that the model efficiency can be improved when deactivating DFI at inference. Additionally, integrating LoRA layers to fine-tune the foundation features is also helpful, leading to a 3.96% improvement in mIoU with marginal additional parameters.
In Fig. 8, we show the qualitative ablation results to visually examine the effects of DFI and LoRA. The segmentation results indicate that DFI is particularly effective at locating tiny and slender parts, while LoRA further enhances performance with refined boundaries and more complete segmentation maps. More ablation studies are provided in the supplementary.